OpenStudy (anonymous):

please help I need to graph the equation using the x and y intercepts x/7-y/5=1

OpenStudy (anonymous):

Do you know what a y-intercept is?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

It is a point that solves the equation where x = 0. The x-intercept is the opposite.

OpenStudy (anonymous):

now i'm comfused

OpenStudy (anonymous):

To find the y-intercept, plug 0 in for x and solve for y. You will come up with a point that is (0,y)

OpenStudy (anonymous):

The intercept on a certain axis is a point where the graph crosses it, two intercepts define the graph of a linear equation. So the value on the other axis is = 0; for y-int, x=0 and for x-int, y=0 First task is to get the equation in intercept form, meaning that the variables should be on different sides of the equation to get to the form: y = mx + b When x = 0, b represents the y-int. Now, getting the equation in the desired form: \[\frac{ x }{ 7 }-\frac{ y }{ 5 }=1\] Subtract x/7: \[-\frac{ y }{ 5 }=1-\frac{ x }{ 7 }\]Multiply by (-5): \[y=\frac{ 5x }{ 7 } - 5\] So, for the y-int, x=0, then: \[y=\frac{ 5*0 }{ 7 } - 5= \frac{ 0 }{ 7 } -5=-5\] Y-int is defined (0,-5) For X-int y=0: \[0=\frac{ 5x }{ 7 } - 5\] Add 5 to both sides: \[\frac{ 5x }{ 7 } =5\] Multiply by 7: \[5x=35\]Divide by 5: \[x =7\] Now, X-int is also defined (7, 0), now place the 2 points which define the line graph and connect them, you're done!  And, that's it sort of, the drawing is sluggish, but you should get it, cheers